The lab manager at a university needs to mix 10% ethanol solution with a 60% ethanol solution to obtain 50 milliliters of a 50% ethanol solution. How milliliters of the 10% solution and of the 60% solution should be mixed?
Answer by stanbon(75887)

The lab manager at a university needs to mix 10% ethanol solution with a 60% ethanol solution to obtain 50 milliliters of a 50% ethanol solution. How milliliters of the 10% solution and of the 60% solution should be mixed?
--------
Equation:
eth + eth = eth
0.10x + 0.60(50-x) = 0.50*50
--------
Multiply thru by 100 to get:
10x + 60*50 - 60x = 50*50
-----
-50x = -10*50
-----
x = 10 ml (amt. of 10% solution to add)
50-x = 40 ml (amt. of 60% solution needed)
=============
Cheers,
Stan H.
===============